Circular No. 7563

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001I IN UGC 2836 S. Beckmann and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova (mag ∼ 17.2) on an unfiltered image taken on Jan. 15.2 UT. The new object was confirmed on earlier images taken on 2000 Dec. 23.2 (mag ∼ 18.4) and 28.3 UT (mag ∼ 17.7). SN 2001I is located at α = 3^{h}43^{m}57^{s}.28, δ = +39^{o}17′39′′.4 (equinox 2000.0), which is 4′′.1 east and 2′′.5 south of the nucleus of UGC 2836. A KAIT image of the same field taken on 2000 Dec. 3.3 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVAE 2001B, 2001C, 2001F, AND 2001H T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra of SN 2001B (cf. IAUC 7555), obtained by M. Calkins on Jan. 14.18 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be a type-Ia supernova near maximum light. The spectrum consists of a blue continuum with well-defined Si II and Ca II features. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 1556 km/s for the host galaxy, the supernova expansion velocity is 7400 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 9 ± 2 days after maximum light.

Matheson et al. add that spectra of SN 2001C (cf. IAUC 7555), obtained by Calkins (as above) on Jan. 14.20 UT, show it to be a type-Ia supernova several days past maximum light. The spectrum consists of a continuum with well-defined Si II and Ca II features. Adopting the NED recession velocity of 3280 km/s for PGC 19975, the supernova expansion velocity is 10 200 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 13 ± 2 days after maximum light.

Matheson et al. also report that spectra of SN 2001F (cf. IAUC 7558), obtained by Calkins on Jan. 14.53 UT, show it to be a type-Ia supernova several weeks past maximum light. The spectrum consists of nebular emission lines of iron and cobalt. They also note that spectra of SN 2001H (cf. IAUC 7560), obtained by Calkins on Jan. 14.16 UT, show it to be a type-II supernova before maximum light. The spectrum consists of an almost-featureless blue continuum with incipient P-Cyg lines of hydrogen.

2001 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green